*T: A Multithreaded Massively Parallel Architecture

Computation Structures Group Memo 325-2 March 5, 1992

(Replaces CSG Memos 325 (Jan 1991) and 325-1 (Nov 1991))

Rishiyur S. Nikhil† Gregory M. Papadopoulos‡ Arvind‡

† Digital Equipment Corporation, Cambridge Research Laboratory ‡ MIT Laboratory for Computer Science

In Proc. 19th. Intl. Symp. on Computer Architecture, May 19-21, 1992, Queensland, Australia

\begin{matrix}c \\ ◯ \end{matrix} ACM 1992

This report describes research done at Digital Equipment Corporation’s Cambridge Research Laboratory and at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the work at MIT is provided in part by the Advanced Research Projects Agency of the Department of Defense under Office of Naval Research contract N00014-89-J-1988.